OpenStudy (anonymous):

I am going to buy exotic fruits. Dragonfruit costs x-4 dollars. Starfruit is five dollars less expensive than rambutan. Rambutan costs 2x dollars more than dragonfruit. How much does it cost to buy one rambutan, two starfruit, and three dragonfruit? Your answer will be an expression that depends on x.

OpenStudy (anonymous):

Dragonfruit = x-4 Rambutan = 2x+(x-4) =3x-4 Starfruit = 3x-4-5 = 3x-9

OpenStudy (anonymous):

now just combine the number of fruits you need and make one equation out of it